OpenStudy (anonymous):

also how can i show that 1/2(1-x^2)^-1/2(-2x)=-x divided by (1-x^2)

OpenStudy (anonymous):

\[1/2(1-x ^{2})^{-1/2}(-2x)=-x/(1-x ^{2})\]

OpenStudy (anonymous):

or\[1/2(1-x ^{2})^{-1/2}(-2x)=-x/\sqrt{1-x ^{2}}\]

OpenStudy (anonymous):

 Iwanbahyudin is correct.